# CHASE: 3D-Consistent Human Avatars with Sparse Inputs via Gaussian Splatting and Contrastive Learning

### v2

@@ Introduction
3D Geometry Contrastive learning
Inspired by the success of contrastive learning in image processing and static point cloud analysis, we advocate for adopting 3D geometry contrastive learning to ensure 3D consistency during animation. To this end, we treat the 3D Gaussians as a 3D point cloud and use DGCNN (Wang et al. 2019) as the feature extractor. The point cloud feature extractor processes the positions of the 3D Gaussians in the observation space ${\mathcal{G}_{o}}$, the adjusted Gaussians ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the selected pose $p_{i}^{a}$, and outputs their features, creating intermediate graph features to capture global geometric information better. The feature vectors are projected into an invariant space. We denote the projected features of ${\mathcal{G}_{o}}$, ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$ as $f_{o}$, $f_{a}$, and $f_{o}^{\prime}$, respectively.
In the invariant space, we aim to maximize the similarity between $f_{a}$ and $f_{o}^{\prime}$, denoted as $\text{D}_{positive}$, and minimize the similarity between $f_{a}$ and $f_{o}$, denoted as $\text{D}_{negative}$. Therefore, we compute the 3D geometry contrastive loss $\mathcal{L}_{contrastive}$ as:
@@
Optimization
@@
Following (Qian et al. 2024), we use the inverse rigid transformation to canonicalize the viewing direction: $\hat{d}=T_{1:3,1:3}^{-1}d$, where $T$ and $d$ is the forward transformation matrix defined in LBS and viewing direction, respectively. Theoretically, canonicalizing viewing direction also promotes consistency of the specular component of canonical 3D Gaussians under rigid transformations.
@@
Following (Qian et al. 2024), SMPL (Loper et al. 2015) parameter fittings from images can be inaccurate. We additionally optimize the per-sequence shape parameter and per-frame translation, global rotation, and local joint rotations.
@@
Our full loss function consists of several components: an RGB loss $\mathcal{L}_{rgb}$, a mask loss $\mathcal{L}_{mask}$, and a perceptual similarity (LPIPS) loss $\mathcal{L}_{LPIPS}$. We compute these losses on both images rendered from the deformed Gaussians $\mathcal{G}_{o}$ and the adjusted Gaussians $\mathcal{G}_{a}$ with their corresponding ground truth images. Additionally, we include a skinning weight regularization loss $\mathcal{L}_{skin}$, as well as isometric regularization losses for both position and covariance, $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$, respectively, following (Qian et al. 2024). We also incorporate a 3D geometry contrastive loss $\mathcal{L}_{contrastive}$:
@@
Figure 3: Qualitative Comparison on ZJU-MoCap (Peng et al. 2021c). We present results for full and sparse inputs (5% of the full inputs) on the ZJU-MoCap dataset. Results show that our CHASE can produce realistic details with both full and sparse inputs, while other approaches struggle to generate smooth details.
Table 1: Quantitative Results on ZJU-MoCap (Peng et al. 2021c). CHASE achieves state-of-the-art performance across every method. The best and the second best results are denoted by pink and yellow. Frames per second (FPS) is measured on an RTX 3090. We train our model on the dataset that includes only 5% of the origin data for fair quantitative comparison. The metrics are reported in the last four rows of the table. LPIPS = LPIPS $\times$ 1000.
@@
NeuralBody (Peng et al. 2021c) 29.07 0.962 52.29 1.5
Ani-NeRF (Peng et al. 2021a) 29.17 0.961 51.98 1.1
HumanNeRF (Weng et al. 2022) 30.24 0.968 31.73 0.3
MonoHuman (Yu et al. 2023) 29.38 0.964 37.51 0.1
3DGS-Avatar (Qian et al. 2024) 30.62 0.965 30.28 50
GauHuman (Hu et al. 2024b) 30.79 0.960 32.73 180
GoMAvatar (Wen et al. 2024) 30.37 0.969 32.53 43
Ours 30.81 0.970 27.48 50
3DGS-Avatar* (Qian et al. 2024) 29.98 0.957 40.01 50
GauHuman* (Hu et al. 2024b) 30.35 0.957 35.68 180
GoMAvatar* (Wen et al. 2024) 30.01 0.958 42.88 43
Ours* 30.48 0.969 29.94 50
Table 2: Quantitative Results on H36M (Ionescu et al. 2013). Our CHASE outperforms current SOTA methods in both full and sparse settings
@@
NARF (Noguchi et al. 2021) 23.00 0.898 22.27 0.881
NeuralBody (Peng et al. 2021c) 22.89 0.896 23.09 0.891
Ani-NeRF (Peng et al. 2021a) 23.00 0.890 22.55 0.880
ARAH (Wang et al. 2022b) 24.79 0.918 23.42 0.896
3DGS-Avatar (Qian et al. 2024) 32.89 0.982 32.50 0.983
@@
Experiment
Dataset
ZJU-MoCap (Peng et al. 2021c). This dataset features multi-view videos captured by 21 cameras, with human poses recorded using a marker-less motion capture system. For our experiments, we selected six sequences (377, 386, 387, 392, 393, 394). Following the protocol established by HumanNeRF (Weng et al. 2022) and 3DGS-Avatar (Qian et al. 2024), we use a single camera for training and the remaining cameras for evaluation. The foreground masks, camera, and SMPL parameters provided by the data set are used for evaluation purposes. We use 5% of the dataset to simulate sparse inputs.
H36M (Ionescu et al. 2013). H36M is another widely used dataset for human avatar research, comprising multi-view videos from four cameras and human poses captured via a marker-based motion capture system. It features multiple subjects performing a variety of complex actions. We conducted experiments on sequences from subjects S1, S5, S6, S7, S8, S9, and S11, selecting representative actions and dividing the videos into training and test frames. Adhering to the protocol set by ARAH (Wang et al. 2022b), we use three cameras, [54138969, 55011271, 58860488],for training and the remaining camera, [60457274],for testing, and follow their preprocessing steps. We use the SMPL parameters and foreground humans following (Peng et al. 2021a).
Comparison with State-of-the-art Methods
We compare our CHASE with various SOTA methods for human avatars, including NerF-based methods such as NeuralBody (Peng et al. 2021c), Ani-NeRF (Peng et al. 2021a), HumanNeRF (Weng et al. 2022), and MonoHuman (Yu et al. 2023), and 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024), GauHuman (Hu et al. 2024b), and GoMAvatar (Wen et al. 2024) under monocular setup on ZJU-MoCap (Peng et al. 2021c). The quantitative results are shown in Tab. 1. Overall, our proposed CHASE achieves the best performance in terms of PSNR, SSIM, and LPIPS with both full and sparse inputs. Our CHASE shows only a small performance drop when using only 5% of the data and significantly outperforms SOTA methods on LPIPS. It is evidence that our method successfully maintains 3D consistency even with sparse inputs.
Qualitative comparisons on novel view synthesis are shown in Fig. 3. We observe that our method preserves more details compared to other SOTA methods. They often struggle to maintain 3D consistency and deliver suboptimal detail reconstruction in human avatar modeling, particularly when only sparse inputs are available.
For H36M (Ionescu et al. 2013), we report the quantitative results against NerF-based methods such as NARF (Noguchi et al. 2021), NeuralBody (Peng et al. 2021c), Ani-NeRF (Peng et al. 2021a) and ARAH (Wang et al. 2022b), and 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024) in Tab. 2. Our CHASE significantly outperforms these methods with sparse inputs, showing that our CHASE generalizes well to novel poses with sparse inputs and reconstructs human avatars with better appearance and geometry detail. For qualitative comparisons on novel pose synthesis, please refer to the Supp.Mat.
Table 3: Ablation Study on ZJU-MoCap (Peng et al. 2021c). We both show the result from full input (top group) and sparse input (bottom group).
@@
Full model 30.41 0.969 29.94 50
Figure 4: Ablation Study on Dynamic Avatar Adjustment, which enhances multi-view consistency, hence improving rendering quality.
@@
Ablation Study
In this section, we conduct ablation experiments using the ZJU-MoCap (Peng et al. 2021c) dataset with both full inputs and sparse inputs to evaluate the effectiveness of our proposed modules. We also conduct experiments to evaluate different backbones in 3D geometry contrastive learning. Notably, our CHASE maintains 3D consistency effectively without increasing any extra inference time.
Non-rigid Deformation. Non-rigid Deformation is designed for complex cloth deformation. As shown in Tab. 3, non-rigid deformation is required to achieve optimal performance, demonstrating non-rigid regions are well rendered.
Dynamic Avatar Adjustment. As shown in Tab.3, incorporating Dynamic Avatar Adjustment (DAA) results in our full model outperforming the baseline in terms of LPIPS. We argue that LPIPS is particularly informative compared to other metrics because it addresses the challenge of accurately reproducing ground-truth appearance from novel views in a monocular setting(Qian et al. 2024; Yang et al. 2024). Fig. 4 illustrates that Dynamic Avatar Adjustment serves as an effective 2D image supervision technique for 3D human body modeling, enhancing 3D consistency and reducing artifacts while improving multi-view consistency.
3D geometry contrastive learning. Our core idea is that cross-3D-modal contrastive learning can facilitate communication between 3D models for obtaining powerful representations. To verify this, we further do ablation studies to show qualitative comparisons in Tab. 3 and Fig. 5. We can find the full model (w/ contrastive) preserves finer details and provides a more realistic and detailed reconstruction of clothing and other complex surfaces, demonstrating that 3D Geometry Contrastive Learning enhances 3D consistency.
Backbone for 3D contrastive learning. In Tab. 3, We show the ablation study on different backbones, including PointNet (Qi et al. 2017) (pointnet) and DGCNN (Wang et al. 2019) (full model). This indicates that DGCNN captures local geometric details better by building a dynamic graph structure, while PointNet relies on global feature learning, which may cause some local information to be missing.
Conclusion
@@
Limitations. 1). CHASE lacks the capability to extract 3D meshes. Developing a method to extract meshes from 3D Gaussians is an important direction for future research. 2). Our CHASE need to find similar pose from dataset. When it comes to monocular videos featuring large-scale movements and diverse actions, it is hard to find pairs that qualify the requirements, which may impact performance.

### v3

@@ 4.2 Dynamic Avatar Adjustment
4.3 3D Geometry Contrastive learning
Inspired by the success of contrastive learning in 2D image processing and static point cloud analysis, we advocate for adopting 3D geometry contrastive learning to ensure 3D consistency of avatars. We treat the 3D Gaussians as a 3D point cloud and use DGCNN [36] as the feature extractor. DGCNN is typically trained on general point cloud datasets, which helps it learn geometric structures. The point cloud feature extractor processes the positions of the 3D Gaussians in the observation space ${\mathcal{G}_{o}}$, the adjusted Gaussians ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the selected pose $p_{i}^{a}$, and outputs their features, creating intermediate graph features to capture global geometric information better. The feature vectors are projected into an invariant space. We denote the projected features of ${\mathcal{G}_{o}}$, ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$ as $f_{o}$, $f_{a}$, and $f_{o}^{\prime}$, respectively.
In the invariant space, we aim to maximize the similarity between $f_{a}$ and $f_{o}^{\prime}$, denoted as $\text{D}_{positive}$, and minimize the similarity between $f_{a}$ and $f_{o}$, denoted as $\text{D}_{negative}$. This feature-level contrastive learning encourages the model to capture pose-aware knowledge, enabling it to differentiate between subtle pose variations while maintaining geometric consistency. Therefore, we compute the 3D geometry contrastive loss $\mathcal{L}_{contrastive}$ as:
@@
4.4 Optimization
We begin by randomly sampling 50k points from the surface of the SMPL mesh to initialize 3D Gaussians in the canonical space.
@@ 4.4 Optimization
Following [29], we use the inverse rigid transformation to canonicalize the viewing direction: $\hat{d}=T_{1:3,1:3}^{-1}d$, where $T$ and $d$ is the forward transformation matrix defined in LBS and viewing direction, respectively. Theoretically, canonicalizing viewing direction also promotes consistency of the specular component of canonical 3D Gaussians under rigid transformations.
@@
Following [29], SMPL [19] parameter fittings from images can be inaccurate. We additionally optimize the per-sequence shape parameter and per-frame translation, global rotation, and local joint rotations.
@@
Our full loss function consists of several components: an RGB loss $\mathcal{L}_{rgb}$, a mask loss $\mathcal{L}_{mask}$, and a perceptual similarity (LPIPS) loss $\mathcal{L}_{LPIPS}$. We compute these losses on both images rendered from the deformed Gaussians $\mathcal{G}_{o}$ and the adjusted Gaussians $\mathcal{G}_{a}$ with their corresponding ground truth images. Additionally, we include a skinning weight regularization loss $\mathcal{L}_{skin}$, as well as isometric regularization losses for both position and covariance, $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$, following [29]. We also incorporate a 3D geometry contrastive loss $\mathcal{L}_{contrastive}$:
@@
Figure 4: Qualitative Comparison on ZJU-MoCap [26]. We present results for full and sparse inputs (5% of the full inputs) on the ZJU-MoCap dataset. Results show that our CHASE can produce realistic details with both full and sparse inputs, while other approaches struggle to generate smooth details.
Figure 5: Qualitative Comparison on H36M [10] with sparse inputs. We demonstrate that our method effectively produces realistic details for novel pose in both rendered images and geometry, whereas other approach struggles to achieve smooth details.
Table 1: Quantitative Results on ZJU-MoCap [26]. CHASE achieves state-of-the-art performance across every method. The best and the second best results are denoted by pink and yellow. Frames per second (FPS) is measured on an RTX 3090. We train our model on the dataset that includes only 5% of the origin data for fair quantitative comparison. The metrics are reported in the last four rows of the table. LPIPS = LPIPS $\times$ 1000.
@@
NeuralBody [26] 52.29 29.07 0.962 1.5
Ani-NeRF [25] 51.98 29.17 0.961 1.1
HumanNeRF [38] 31.73 30.24 0.968 0.3
MonoHuman [44] 37.51 29.38 0.964 0.1
DVA [31] 37.74 29.45 0.956 17
InstantAvatar [11] 64.41 29.73 0.938 4.2
3DGS-Avatar [29] 30.28 30.62 0.965 50
GauHuman [8] 32.73 30.79 0.960 180
GoMAvatar [37] 32.53 30.37 0.969 43
Ours 27.48 30.81 0.970 50
3DGS-Avatar* [29] 40.01 29.98 0.957 50
GauHuman* [8] 35.68 30.35 0.957 180
GoMAvatar* [37] 42.88 30.01 0.958 43
Ours* 29.94 30.48 0.969 50
5 Experiment
5.1 Dataset
ZJU-MoCap [26]. This dataset features multi-view videos captured by 21 cameras, with human poses recorded using a marker-less motion capture system. For our experiments, we selected six sequences (377, 386, 387, 392, 393, 394). Following the protocol established by HumanNeRF [38] and 3DGS-Avatar [29], we use a single camera for training and the remaining cameras for evaluation. The foreground masks, camera, and SMPL parameters provided by the data set are used for evaluation purposes. We simulate sparse inputs by sampling every 20th frame from the video, between start_frame and end_frame, which corresponds to using 5% of the total frames. We also specify in the Supp.Mat which images are selected for training.
H36M [10]. H36M is another widely used dataset for human avatar research, comprising multi-view videos from four cameras and human poses captured via a marker-based motion capture system. We conducted experiments on sequences from subjects S1, S5, S6, S7, S8, S9, and S11, selecting representative actions and dividing the videos into training and test frames. Adhering to the protocol set by ARAH [35], we use three cameras, [54138969, 55011271, 58860488], for training and the remaining camera, [60457274], for testing, and follow their preprocessing steps. We use the SMPL parameters and foreground humans following [25]. We apply the same approach as mentioned above to simulate sparse inputs in this dataset.
Table 2: Quantitative Results on H36M [10]. Our CHASE outperforms current SOTA methods in both full and sparse settings
@@ 5.1 Dataset
NARF [24] 23.00 0.898 22.27 0.881
NeuralBody [26] 22.89 0.896 23.09 0.891
Ani-NeRF [25] 23.00 0.890 22.55 0.880
ARAH [35] 24.79 0.918 23.42 0.896
3DGS-Avatar [29] 32.89 0.982 32.50 0.983
@@
5.2 Comparison with State-of-the-art Methods
We compare our CHASE with various SOTA methods for human avatars, including NerF-based methods such as NeuralBody [26], Ani-NeRF [25], HumanNeRF [38], and MonoHuman [44], and 3DGS-based methods such as 3DGS-Avatar [29], GauHuman [8], and GoMAvatar [37] under monocular setup on ZJU-MoCap [26]. The quantitative results are shown in Tab. 1. Overall, our proposed CHASE achieves the best performance in terms of PSNR, SSIM, and LPIPS with both full and sparse inputs. Notably, our CHASE shows only a small performance drop when using only 5% of the data, especially in LPIPS which is more informative than the other two metrics in our setting [29]. In fact, it declines less compared to other methods and even surpasses their performance with 100% of the data. It is evidence that our method successfully maintains 3D consistency even with sparse inputs.
Qualitative comparisons on novel view synthesis are shown in Fig. 4. We observe that our method preserves more details compared to other SOTA methods. They often struggle to maintain 3D consistency and deliver suboptimal detail reconstruction in human avatar modeling, particularly when only sparse inputs are available. Please see our project website and supplementary material for more visualization.
For H36M [10], we report the quantitative results against NeRF-based methods such as NARF [24], NeuralBody [26], Ani-NeRF [25] and ARAH [35], and 3DGS-based methods such as 3DGS-Avatar [29] in Tab. 2. Our CHASE significantly outperforms these methods with sparse inputs, showing that our CHASE generalizes well to novel poses with sparse inputs and reconstructs human avatars with better appearance and geometry detail. For qualitative comparisons on novel pose synthesis, as shown in Fig. 5, our method generalizes well to novel pose with just sparse inputs (only 5% of origin data) and reconstruct human avatars with better appearance and geometry detail.
Table 3: Ablation Study on ZJU-MoCap [26]. We both show the result from full input (top group) and sparse input (bottom group).
@@ 5.2 Comparison with State-of-the-art Methods
Top-3 30.44 0.969 29.98 50
Top-5 30.43 0.968 30.19 50
Full model 30.48 0.969 29.94 50
Figure 6: Ablation Study on DAA, which enhances multi-view 3D consistency, hence improving the overall rendering quality.
@@
5.3 Ablation Study
In this section, we conduct ablation experiments using the ZJU-MoCap [26] dataset with both full inputs and sparse inputs to evaluate the effectiveness of our proposed modules. We also conduct experiments to evaluate different backbones in 3D geometry contrastive learning. Notably, our CHASE maintains 3D consistency effectively without increasing any extra inference time.
Non-rigid deformation. Non-rigid Deformation is designed for complex deformations, such as cloth bending and stretching. As shown in Tab. 3, non-rigid deformation is required to achieve optimal performance, demonstrating non-rigid regions are well rendered.
Dynamic Avatar Adjustment. As shown in Tab.3, incorporating DAA results in our full model outperforming the baseline in terms of LPIPS which is particularly informative compared to other metrics in our setting [29, 40]. Fig. 6 illustrates that DAA serves as an effective 2D image supervision for 3D human body modeling, enhancing 3D consistency and reducing artifacts while improving multi-view consistency. Additionally, DAA reduces artifacts and inaccuracies in the geometry caused by sparse inputs, further enhancing the robustness and visual realism of the model in practical applications.
Similar (pose/image) selection. Our method selects a similar (pose/image) to supervise the generated avatar as described in Section 4.2. Ablation studies in Tab. 3 show that even when selecting less similar (poses/images), our approach significantly improves performance. Top-n refers to choosing the n-th most similar pose/image. These results highlight CHASE ’s effectiveness in handling limited data or irregular large-scale human movements.
3D geometry contrastive learning. Our core idea is that cross-3D-modal contrastive learning can facilitate communication between 3D models for obtaining powerful representations. To verify this, we further do ablation studies to show qualitative comparisons in Tab. 3 and Fig. 7. We can find the full model (w/ contrastive) preserves finer details and provides a more realistic and detailed reconstruction of clothing, demonstrating that 3D geometry contrastive learning enhances 3D consistency.
Backbone for 3D contrastive learning. In Tab. 3, we show the ablation study on different backbones, including PointNet [27] (pointnet) and DGCNN [36] (full model). This indicates that DGCNN captures local geometric details better by building a dynamic graph structure, while PointNet relies on global feature learning, which may cause some local information to be missing.
6 Conclusion
@@
Limitations. 1). CHASE lacks the capability to extract 3D meshes. Developing a method to extract meshes from 3D Gaussians is an important direction for future research. 2). CHASE performs less effectively in reconstructing humans with complex clothing, as similar poses may still exhibit substantial differences in non-rigid deformations. This remains a challenge we plan to address in future work.